# Adaptive Bounding Box Uncertainties via Two-Step Conformal Prediction

### v1

@@ 6 Experiments
For our experiments we rely on pre-trained object detection models from detectron2 [69], primarily based on a Faster R-CNN architecture and trained on COCO [32]. We consider datasets COCO (validation set), Cityscapes [13] and BDD100k [72], which contain 2D bounding box annotations and are split into calibration and test data respectively. We run our two-step conformal approach for a variety of classes, and focus our results on a coherent set of COCO classes to which class labels across all datasets can be mapped to: person, bicycle, motorcycle, car, bus and truck (see subsection 0.B.3).
Since an image may contain multiple objects, we require a pairing between an object’s true and predicted bounding boxes. Similarly to [14, 1] we perform Hungarian matching [25] based on an intersection-over-union (IoU) threshold of 0.5. Throughout, we set $\alpha_{L}=0.01,\alpha_{B}=0.1$ for a nominal box coverage of $(1-\alpha_{L})(1-\alpha_{B})\approx 90\%$, and employ $\max$-rank [58] for multiple testing correction. Results are collected over repeated trials of random calibration and test splits. Additional results, including different coverage level combinations $(\alpha_{L},\alpha_{B})$, can be found in Appendix 0.C.
Figure 3: Top: Empirical coverage levels marginally across all objects (All) and across objects from selected classes for the three conformal bounding box methods on the BDD100k dataset. Target coverage is achieved both marginally and for individual classes. Bottom: Coverage levels are stratified by object size (Small, Medium, Large), showing that Box-CQR and in particular Box-Ens provide a more balanced empirical coverage across sizes. However, this comes at the cost of slightly larger intervals, as seen when comparing $MPIW$. We also visualize target coverage (Refer to caption) and the marginal coverage distribution (Refer to caption). Displayed densities are results obtained over 1000 trials.
@@ 6.1 Metrics
Let us consider a test dataset $\mathcal{D}_{test}=\{(X_{j},Y_{j})\}_{j=n+1}^{n+n_{t}}$ of size $n_{t}$. We validate our approach by assessing our key desiderata using relevant CP metrics as described below, which jointly capture desired notions of ‘reliable’ uncertainty [2].
Validity. We ensure that our CP procedure satisfies nominal coverage guarantees by verifying its empirical coverage, which we define in generality as
@@
where $\mathbbm{1}[\cdot]$ is the indicator function which takes the form $\mathbbm{1}[l_{j}\in\hat{C}_{L}(X_{j})]$ for label prediction sets and $\mathbbm{1}[\bigcap_{k=1}^{m}(c_{j}^{k}\in\hat{C}^{k}_{B}(X_{j}))]$ for bounding box PIs. Note that $Cov$ is a random quantity parametrized by an empirical coverage distribution and may deviate from nominal coverage levels based on factors such as calibration set size [60].
Adaptivity. To verify if target coverage is achieved by compensating undercoverage on some objects with overcoverage on others, similarly to [5, 49] we verify empirical coverage across different strata, namely object sizes. We follow the COCO detection challenge [^3] https://cocodataset.org/#detection-eval and stratify across three bounding box sizes by surface area: small ($Cov_{S}$, area $\leq 32^{2}$), medium ($Cov_{M}$, area $\in(32^{2},96^{2}]$) and large ($Cov_{L}$, area $>96^{2}$).
Efficiency. Obtained conformal prediction sets and intervals are desired to be as small as possible while still satisfying target coverage (i.e., remaining valid). We define mean set size for the label prediction sets and mean prediction interval width ($MPIW$) for the bounding box PIs as
@@
Predictive performance. We also follow standard practice and validate model performance using object detection-specific metrics from the COCO challenge, i.e., average precision across multiple IoU thresholds (see Appendix 0.C).
@@
Method $MPIW$ $Cov$ $MPIW$ $Cov$
DeepEns [27] $12.31\pm 0.47$ $0.21\pm 0.01$ $74.15\pm 2.01$ $0.49\pm 0.01$
GaussianYOLO [12] $7.00\pm 0.14$ $0.08\pm 0.01$ $87.07\pm 4.25$ $0.35\pm 0.01$
Andéol et al. (Best) Not Applicable $87.62\pm 1.79$ $0.91\pm 0.01$
Box-Std (Ours) $55.47\pm 2.97$ $0.88\pm 0.02$ $85.42\pm 1.99$ $0.88\pm 0.02$
Table 2: We compare our simplest method Box-Std to Andéol et al.’s best results (see subsection 0.B.5) as well as deep ensembles (DeepEns) and GaussianYOLO, two popular UQ approaches. The former is valid (green) but only designed for one-sided intervals, while the latter two heavily undercover (red), highlighting the unreliability of obtained uncertainties due to missing guarantees. Results are for COCO across the set of selected classes and 100 trials ($\alpha_{B}=0.1$).
@@ 6.2 Comparison of bounding box methods
Empirical coverage levels stratified by class labels as well as object sizes for the three proposed bounding box methods are displayed in Figure 3 for the BDD100k dataset. We observe that the target coverage of 90% is satisfied per class, validating the class-conditional guarantee provided in Equation 3. The displayed magnitudes of variation are explained by the differences in available calibration samples per class (see subsection 0.A.6 and Table 3).
We further observe that the fixed-width PI construction of Box-Std may be wide enough to cover most small objects, but will fail to account for the magnitude of large ones, with significant undercoverage. In contrast, the adaptive nature of Box-CQR and in particular Box-Ens via its uncertainty-dependent scaling can account for varying magnitudes, achieving higher coverage for large objects at a slight loss in efficiency due to the resulting wider intervals. Whilst coverage across small objects reduces somewhat, it now intuitively aligns with observed prediction difficulty (see Table 6). That is, objects which are more challenging to predict exhibit a higher variation and chance of miscoverage. The improved balance in coverage across object sizes is a purely empirical benefit of our adaptive designs – the employed CP procedure only aims to guarantee target coverage per class and does not condition on object size.
Baseline comparisons. We validate results by comparing to Andéol et al. [1, 14], modifying our own two-sided interval methods to produce one-sided PIs, and evaluating efficiency via $MPIW$ and their box stretch metric (see subsection 0.B.5 for details). Table 2 shows that we achieve marginally tighter intervals even in their own, more restricted setting, while remaining valid. We additionally evaluate generated uncertainties via deep ensembles (DeepEns) [27] and GaussianYOLO [12], two popular UQ approaches for object detection (see section 3). Results confirm the unreliability of produced uncertainties due to their lack of guarantees, as seen by severe undercoverage in both settings. Finally, a comparison of the $\max$-rank correction to Bonferroni in Table 5 asserts that substantially tighter PIs can be obtained, while overcoverage tendencies are surpressed.
[…]
@@ 6.3 Results for the two-step approach
We compare our proposed two-step approach, i.e., employing any conformal box method with a preceeding conformal step for the class labels via ClassThr, against the proposed baselines for label prediction sets in Figure 4. Each label set method’s nominal guarantee and expected efficiency is displayed in Table 1.
In line with our expectations, only approaches using ClassThr or Full (i.e., the full label set) consistently achieve target coverage for both class labels and subsequent bounding box PIs across all three datasets. Notably, while Full results in overly conservative box PIs due to its set construction, ClassThr provides surprisingly efficient label sets (with a mean set size of at most four) which translate to actionably tight box intervals. Differences exist in the efficiency of the different conformal box methods, with Box-Ens performing notably better on BDD100k. Stratifying results by object (mis-)classification supports anticipated behaviour, in that wrongly classified objects tend to exhibit slightly higher uncertainty (see Figure 8).
@@
Oracle relies on knowing the correct quantile and thus does not require generating a label set. While it provides nominal guarantees assuming correct label prediction and empirically satisfies them with high efficiency (i.e., narrow PIs), the condition severely limits its practicality. Top consistently undercovers the true label, and may tend to slightly undercover boxes as sample sizes increase. Naive is able to consistently maintain box coverage even though label coverage is violated, with surprisingly narrow PIs. However, additional experiments in subsection 0.C.1 showcase its sensitivity to model miscalibration, yielding it less robust than ClassThr. Yet, it is interesting that methods such as Top and Naive perform quite well empirically even under void nominal guarantees. Among others, this relates back to comments made on dominating label quantiles in subsection 5.1.
@@
Discussion and method choice. We end by discussing the choices in selecting a suitable conformal two-step approach a practitioner may want to consider. In terms of label set methods, we note the following: if the model classifier records a high accuracy and there exists a way to externally validate predicted class labels, then Top may be a highly efficient choice. If the model is strongly calibrated and only empirical assurances are sufficient, then Naive may be a suitable selection. However, if maximal safety assurances with full nominal and empirical guarantees are desired, then ClassThr is the only safe choice. We highlight that improvement potential remains: obtained PIs using ClassThr tend to overcover, presumably due to our conservative quantile selection strategy. It may be indeed possible to obtain the same set of safety assurances with even higher efficiency using a different strategy. For instance, one could consider a weighted quantile construction on the basis of the classifier’s confusion matrix.
@@
Regarding the choice of conformal box method, we observe that Box-Std will be the most efficient if the object detection task contains only similarly sized objects. However, if objects vary substantially in size, an adaptive approach such as Box-Ens or Box-CQR will be more suitable. One could also consider designing a CP approach which explicitly satisfies target coverage over other reasonable strata beyond classes, such as object sizes or shapes.
@@ 7 Conclusion
We present a novel procedure to quantify predictive uncertainty for multi-object detection. We leverage CP to generate uncertainty intervals with a per-class coverage guarantee for new samples. Our proposed two-step conformal approach provides adaptive bounding box intervals with safety assurances robust to object misclassification. We aim to extend our approach to 3D bounding boxes, object tracking and other detection tasks in future work, focussing on similar practically useful guarantees. Whilst further improvements can be made to achieve narrower intervals, our results are promisingly tight, paving the way for a safer deployment of vision-based systems in scenarios involving decision-making under uncertainty.
@@ 0.A.1 Split conformal prediction
2:Output: Prediction set $\hat{C}(X_{n+1})$ for test sample $(X_{n+1},Y_{n+1})$.
@@
8:For a new test sample $(X_{n+1},Y_{n+1})$, a valid conformal prediction set for $X_{n+1}$ is given by $\hat{C}(X_{n+1})=\{y\in\mathcal{Y}:s(\hat{f}(X_{n+1}),y)\leq\hat{q}\}$. Validity refers to satisfying sample coverage with marginal probability $(1-\alpha)$, i.e., the guarantee in Equation 1. We refer to the relevant literature for the proof.
@@
Rather than a marginal guarantee, we desire the class-conditional guarantee of Equation 2, which is re-interpreted as Equation 4 for our object detection setting outlined in section 2. Thus, we may relate our problem setting to Algorithm 1 as follows:
@@
    A calibration sample $(X_{i},Y_{i})$ consists of an input image $X_{i}$ and an existing object in that image, parametrized as $(c^{1}_{i},\dots,c^{m}_{i},l_{i})$ for an arbitrary amount of coordinates $m$ following subsection 2.3. As we run Algorithm 1 on a per-class basis, the class label $l$ for samples within a class is fixed, and we thus define $Y_{i}:=(c^{1}_{i},\dots,c^{m}_{i})$ as the object’s coordinates.
@@
    Finally, the conformal procedure is performed for each of the $m$ coordinates in parallel, therefore necessating a multiple testing correction as motivated in subsection 4.1 and subsection 0.A.3. Thus, we define the final set of calibration samples for the $k$-th coordinate of an object belonging to class $y$ as
@@
In line with our notation in section 4, the required values to compute a nonconformity score for $(X_{i},c^{k}_{i})\in\mathcal{D}_{cal,y,k}$ are the true $k$-th coordinate $c^{k}_{i}$ and the model’s prediction $\hat{f}_{B}(X_{i})=\hat{c}^{k}_{i}$; the computed conformal quantile for $\mathcal{D}_{cal,y,k}$ is denoted $\hat{q}_{B}^{k}$; and the obtained prediction interval for a test sample’s $k$-th coordinate $(X_{n+1},c^{k}_{n+1})$ is denoted $\hat{C}^{k}_{B}(X_{n+1})$. The miscoverage rate becomes $\alpha_{B}$.

### v2

@@ 6 Experiments
For our experiments we primarily rely on pre-trained object detectors from detectron2 [72], based on a Faster R-CNN architecture and trained on COCO [33]. We consider three datasets: COCO validation, Cityscapes [14] and BDD100k [75], which contain 2D bounding box annotations and are split into appropriate calibration and test sets. We run our two-step conformal procedure for a variety of classes, but focus reported results on a coherent set of object classes which exists across datasets: person, bicycle, motorcycle, car, bus and truck (see § 0.B.3).
Since the images can contain multiple objects, we require a pairing mechanism between true and predicted bounding boxes. Following prior work [15, 1] we perform Hungarian matching [26] based on an intersection-over-union (IoU) threshold of 0.5. Throughout, we set $\alpha_{L}=0.01,\alpha_{B}=0.1$ for a target box coverage of $(1-\alpha_{L})(1-\alpha_{B})\approx 90\%$, and employ $\max$-rank [61] for multiple testing correction. Results are averaged across multiple trials of data splitting. Additional results, including varying combinations of $(\alpha_{L},\alpha_{B})$, are in Appendix 0.C  [^3] Our code is publicly available at https://github.com/alextimans/conformal-od..
Figure 3: Top: Empirical coverage levels marginally across all objects (All) and across objects from selected classes for the three bounding box methods (§ 4) on the BDD100k dataset. Target coverage is achieved both marginally and for individual classes. Bottom: Coverage levels are stratified by object size (Small, Medium, Large), showing that Box-CQR and in particular Box-Ens provide a more balanced empirical coverage across sizes. However, this comes at the cost of slightly larger intervals, as seen when comparing $MPIW$. We also visualize target coverage (Refer to caption) and the marginal coverage distribution (Refer to caption). Displayed densities are results obtained over 1000 trials.
@@ 6.1 Metrics
Our approaches are validated by assessing the key desiderata of CP via relevant metrics described below, which jointly capture the desired notions of ‘reliable’ uncertainty [2]. We denote the test set of size $n_{t}$ as $\mathcal{D}_{test}=\{(X_{j},Y_{j})\}_{j=n+1}^{n+n_{t}}$.
Validity. We assess if nominal coverage guarantees are satisfied by verifying empirical coverage, which we define in generality as
@@
where $\mathbbm{1}[\cdot]$ is the indicator function, of form $\mathbbm{1}[l_{j}\in\hat{C}_{L}(X_{j})]$ for label prediction sets and $\mathbbm{1}[\bigcap_{k=1}^{m}(c_{j}^{k}\in\hat{C}^{k}_{B}(X_{j}))]$ for box intervals. Note that $Cov$ is a random quantity parametrized by an empirical coverage distribution, and will deviate from nominal coverage based on factors such as calibration set size $|\mathcal{D}_{cal}|$ [63].
Adaptivity. To examine if target coverage is satisfied by an imbalance of over- and undercoverage across objects, similarly to [5, 51] we verify empirical coverage by stratification, namely across object sizes. We follow the COCO challenge  [^4] See https://cocodataset.org/#detection-eval. and stratify across three sizes by bounding box surface area: small ($Cov_{S}$, area $\leq 32^{2}$), medium ($Cov_{M}$, area $\in(32^{2},96^{2}]$) and large ($Cov_{L}$, area $>96^{2}$).
Efficiency. Obtained conformal prediction sets and intervals are desired to be as small as possible while still maintaining target coverage (i.e., remaining valid). We define the mean set size for label prediction sets and mean prediction interval width ($MPIW$) for bounding box prediction intervals as
@@
Predictive performance. We also follow standard practice and validate model performance using object detection-specific metrics from the COCO challenge, in particular average precision across multiple IoU thresholds (see Appendix 0.C).
@@
Uncertainty method Object detector $MPIW$ $Cov$ $MPIW$ $Cov$
DeepEns 5×5$\times$5 × Faster R-CNN $12.31\pm 0.47$ $0.21\pm 0.01$ $74.15\pm 2.01$ $0.49\pm 0.01$
GaussianYOLO YOLOv3 $7.00\pm 0.14$ $0.08\pm 0.01$ $87.07\pm 4.25$ $0.35\pm 0.01$
Andéol et al. (Best) Faster R-CNN N/A $87.62\pm 1.79$ $0.91\pm 0.01$
YOLOv3 N/A $107.93\pm 4.85$ $0.92\pm 0.02$
DETR N/A $82.21\pm 1.64$ $0.90\pm 0.01$
Sparse R-CNN N/A $79.35\pm 1.78$ $0.91\pm 0.01$
Box-Std (Ours) Faster R-CNN $55.47\pm 2.97$ $0.88\pm 0.02$ $85.42\pm 1.99$ $0.88\pm 0.02$
YOLOv3 $61.73\pm 3.66$ $0.88\pm 0.02$ $103.12\pm 3.95$ $0.88\pm 0.02$
DETR $45.34\pm 3.33$ $0.88\pm 0.02$ $80.57\pm 1.78$ $0.88\pm 0.01$
Sparse R-CNN $41.92\pm 2.16$ $0.89\pm 0.01$ $77.33\pm 1.72$ $0.89\pm 0.01$
Table 2: We compare our simplest method Box-Std to Andéol et al.’s best results (see § 0.B.5) across different object detectors (Faster R-CNN [72], YOLOv3 [48], DETR [10], Sparse R-CNN [59]) as well as deep ensembles (DeepEns [28]) and GaussianYOLO [13], two popular UQ approaches. The former satisfies coverage but is only designed for one-sided intervals, while the latter can heavily undercover in practice (marked [Uncaptioned image]). Results are for COCO across classes and 100 trials, for target coverage $(1-\alpha_{B})=0.9$. The key difference between various object detectors are the obtained interval widths ($MPIW$), which relate to predictive performance and are smaller for better models.
@@ 6.2 Comparison of bounding box methods
Empirical coverage levels stratified by class labels as well as object sizes for the three proposed bounding box methods are displayed in Fig. 3 for BDD100k. We see that target coverage of 90% is satisfied even per class, validating the class-conditional guarantees provided in Eq. 3. The visible coverage variations are explained by the differences in available calibration samples per class (see § 0.A.6 and Tab. 3). We further observe that the fixed-width intervals of Box-Std may be large enough to cover small objects, but will fail to account for the magnitude of large ones, resulting in significant undercoverage. In contrast, the adaptive nature of Box-CQR and in particular Box-Ens via its scaling factor can better account for varying magnitudes, achieving higher coverage for large objects at a slight loss in efficiency due to a higher $MPIW$. Whilst coverage across small objects reduces somewhat, it now intuitively aligns with observed prediction difficulty (see Tab. 6). That is, objects which are more challenging to predict exhibit a higher variation and chance of miscoverage. We note that the improved coverage balance across object sizes is a purely empirical benefit of our adaptive designs – the employed conformal procedures only aim to guarantee target coverage per class and do not condition on object size.
Baseline comparisons. We further validate our conformal bounding box step by comparing to Andéol et al. [1, 15], modifying our own two-sided interval methods to produce one-sided PIs, and evaluating efficiency via $MPIW$ (see § 0.B.5 for details and their box stretch metric). Tab. 2 demonstrates that we achieve marginally tighter intervals even in their own, more restricted setting, while remaining equally valid. We also evaluate generated uncertainties via deep ensembles (DeepEns) [28] and GaussianYOLO [13], two popular UQ approaches for object detection. Results confirm the unreliability of produced uncertainties due to their lack of guarantees, as seen by severe undercoverage in both settings. Finally, a comparison of the $\max$-rank correction to Bonferroni in § 0.C.2 asserts that substantially tighter PIs can be obtained with our employed correction.
Figure 4: Every combination of conformal label set and bounding box method is evaluated along two axes for COCO (top row), Cityscapes (middle row) and BDD100k (bottom row). On the vertical axis we display efficiency, i.e., mean set size for label sets (left column) and $MPIW$ for box intervals (right column). On the horizontal axis we display empirical coverage levels. We also draw target coverage (Refer to caption) and marginal coverage distributions (Refer to caption). In line with Tab. 1, approaches employing ClassThr or Full consistently achieve both label and box target coverage, at the cost of larger prediction sets/intervals. Results are averaged across classes and 100 trials.
@@ 6.3 Results for the two-step approach
After having benchmarked our bounding box step, we next compare the full two-step approach by adding the preceeding conformal step for class labels via ClassThr, and comparing to the proposed label set baselines in § 5.3. Each method’s nominal guarantee and expected efficiency is displayed in Tab. 1.
@@
In line with expectations, we observe in Fig. 4 that only approaches using ClassThr or Full (the full label set) consistently achieve target coverage for both class labels and box intervals across all three datasets. Notably, while Full results in overly inflated interval widths due to its construction, ClassThr provides surprisingly efficient label sets (with mean set size $\leq 4$) which propagate into reasonably tight box intervals. Differences also exist in the efficiency of the three bounding box methods, with Box-Ens performing notably better on BDD100k. Stratifying results by object (mis-)classification also confirms that wrongly classified objects tend to exhibit higher uncertainty (see Fig. 8). Oracle relies on knowing the correct quantile and thus does not require generating a label set. While it provides nominal guarantees assuming correct label prediction and empirically satisfies them with high efficiency (i.e., small $MPIW$), the condition severely limits its practicality. Top consistently undercovers the true label, and also tends towards undercovering boxes as sample sizes increase. Naive is able to consistently maintain box coverage even though label coverage is violated, with surprisingly tight PIs. However, additional experiments in § 0.C.1 showcase its sensitivity to model miscalibration, yielding it less robust than ClassThr. Yet, it is interesting that methods such as Top and Naive perform quite well empirically, even under void nominal guarantees.
@@
Discussion and practitioner’s choice. We end by discussing the choices in selecting a suitable conformal two-step approach a practitioner may want to consider. In terms of label set methods, we suggest the following: if the model classifier records a high accuracy and there exists a way to externally validate predicted class labels, then Top may be a highly efficient choice. If the model is strongly calibrated and only empirical assurances are sufficient, then Naive may be a suitable selection. However, if strong safety assurances with both nominal and empirical guarantees are desired, then ClassThr is the only safe choice. We highlight that improvement potential remains: obtained PIs using ClassThr tend to overcover, presumably due to our conservative quantile selection strategy. It may be indeed possible to obtain the same set of safety assurances with higher efficiency under a different strategy. For example, one may consider a weighted quantile construction on the basis of the classifier’s confusion matrix.
@@
Regarding the choice of bounding box method, we observe that Box-Std will be the most efficient if the object detection task contains only similarly sized objects. However, if objects vary substantially in size, an adaptive approach such as Box-Ens or Box-CQR will be more suitable. One may also consider designing a conformal approach which explicitly satisfies target coverage across other reasonable strata beyond classes, such as object sizes or shapes. A limiting factor to consider may be the size of available calibration data per partition.
@@ 7 Conclusion
We present a novel procedure to quantify predictive uncertainty for multi-object detection. We leverage CP to generate uncertainty intervals with a per-class coverage guarantee for new samples. Our proposed two-step conformal approach provides adaptive bounding box intervals with safety assurances robust to object misclassification. Addressing similar types of guarantees, the procedure can be extended to 3D bounding boxes, object tracking and other detection tasks in future work. Whilst improvements can be made to achieve even narrower intervals, our results are promisingly tight, paving the way for a safer deployment of vision-based systems in scenarios involving decision-making under uncertainty.
@@ 0.A.1 Split conformal prediction
2:Output: Prediction set/interval $\hat{C}(X_{n+1})$ for test sample $(X_{n+1},Y_{n+1})$.
@@
8:For a new test sample $(X_{n+1},Y_{n+1})$, a valid prediction set/interval for $X_{n+1}$ is given by $\hat{C}(X_{n+1})=\{y\in\mathcal{Y}:s(\hat{f}(X_{n+1}),y)\leq\hat{q}\}$. Validity refers to satisfying sample coverage with marginal probability $(1-\alpha)$, i.e., the guarantee in Eq. 1. We refer to the relevant literature for the proof.
@@
Rather than a marginal guarantee, we desire the class-conditional guarantee of Eq. 2, which is re-interpreted as Eq. 4 for our object detection setting outlined in § 2. Thus, we may relate our problem setting to Algorithm 1 as follows:
@@
    A calibration sample $(X_{i},Y_{i})$ consists of an input image $X_{i}$ and an existing object in that image, parametrized as $(c^{1}_{i},\dots,c^{m}_{i},l_{i})$ for an arbitrary amount of coordinates $m$ following § 2.3. As we run Algorithm 1 on a per-class basis, the class label $l$ for samples within a class is fixed, and we thus define $Y_{i}:=(c^{1}_{i},\dots,c^{m}_{i})$ as the object’s coordinates.
@@
    Finally, the conformal procedure is performed for each of the $m$ coordinates in parallel, therefore necessating a multiple testing correction as motivated in § 4.1 and § 0.A.3. Thus, we define the final set of calibration samples for the $k$-th coordinate of an object belonging to class $y$ as
@@
In line with our notation in § 4, the required values to compute a nonconformity score for $(X_{i},c^{k}_{i})\in\mathcal{D}_{cal,y,k}$ are the true $k$-th coordinate $c^{k}_{i}$ and the model’s prediction $\hat{f}_{B}(X_{i})=\hat{c}^{k}_{i}$; the computed conformal quantile for $\mathcal{D}_{cal,y,k}$ is denoted $\hat{q}_{B}^{k}$; and the obtained prediction interval for a test sample’s $k$-th coordinate $(X_{n+1},c^{k}_{n+1})$ is denoted $\hat{C}^{k}_{B}(X_{n+1})$. The miscoverage rate becomes $\alpha_{B}$.